# Zero-shot Logical Query Reasoning on any Knowledge Graph

### 10 Apr 2024 v1

@@ Abstract
Complex logical query answering (CLQA) in knowledge graphs (KGs) goes beyond simple KG completion and aims at answering compositional queries comprised of multiple projections and logical operations. Existing CLQA methods that learn parameters bound to certain entity or relation vocabularies can only be applied to the graph they are trained on which requires substantial training time before being deployed on a new graph. Here we present UltraQuery, an inductive reasoning model that can zero-shot answer logical queries on any KG. The core idea of UltraQuery is to derive both projections and logical operations as vocabulary-independent functions which generalize to new entities and relations in any KG. With the projection operation initialized from a pre-trained inductive KG reasoning model, UltraQuery can solve CLQA on any KG even if it is only finetuned on a single dataset. Experimenting on 23 datasets, UltraQuery in the zero-shot inference mode shows competitive or better query answering performance than best available baselines and sets a new state of the art on 14 of them. The code is available: https://github.com/DeepGraphLearning/ULTRA
Graph Machine Learning, Knowledge Graphs
@@ 1 Introduction
Complex logical query answering (CLQA) generalizes simple knowledge graph (KG) completion to more complex, compositional queries with logical operators such as intersection $(\wedge)$, union $(\lor)$, and negation $(\lnot)$. Such queries are expressed in a subset of first-order logic (FOL) where existentially quantified $(\exists)$ variables and given constants comprise relation projections (or atoms), and logical operators combine projections into a logical query (graph pattern). A typical example of a logical query (Ren et al., 2023) is presented in Figure 1: $U.\exists V:\texttt{Win}(\texttt{NobelPrize},V)\land\texttt{Citizen}(\texttt{USA},V)\land\texttt{Graduate}(V,U)$ where $\texttt{Win}()$ is a relation projection, NobelPrize is a constant, and $V$ is an existentially quantified variable.
@@
Due to the incompleteness of most KGs, these logical queries cannot be directly solved by graph traversal algorithms. Consequently, CLQA methods have to deal with missing edges when modeling the projection operators. The vast majority of existing CLQA methods (Ren et al., 2023) predict missing edges by learning graph-specific entity and relation embeddings making such approaches transductive and not generalizable to other KGs. A few approaches (Galkin et al., 2022b) are able to generalize query answering to new nodes at inference time but still need a fixed relation vocabulary.
@@
In this work, we focus at the hardest inductive generalization setup where queries and underlying graphs at inference time are completely different from the training graph, i.e., both entities and relations are new. Furthermore, we aim at performing CLQA in the zero-shot setting with one single model. That is, instead of adaptively finetuning a model on each target dataset, we seek to design a unified approach that generalizes to any KG and any complex query at inference time. For example, in Figure 1, the training graph describes academic entities with relations Win, Citizen, Graduate [^1] We assume the presence of respective inverse relations $\texttt{r}^{-1}$. whereas the inference graph describes music entities with relations Band Member and Plays. The query against the inference graph $U:\texttt{BandMember}(\texttt{Dire Straits},U)\wedge\texttt{Plays}^{-1}(\texttt{Guitar},U)$ involves both new entities and relations and, to the best of our knowledge, cannot be tackled by any existing CLQA method that learns a fixed set of entities or relation embeddings from the training graph.
@@
To this end, we leverage the recent progress in foundation models for KG reasoning (Galkin et al., 2023; Gao et al., 2023) and devise UltraQuery, the first CLQA approach that generalizes to logical queries on any arbitrary KG with any entity and relation vocabulary in the zero-shot fashion without relying on any external node or edge features. UltraQuery parameterizes the projection operator by an inductive graph neural network (GNN) and implements non-parametric logical operators with fuzzy logics (van Krieken et al., 2022). The pre-trained projection operator (Galkin et al., 2023) does not learn any graph-specific entity nor relation embeddings thanks to the generalizable meta-graph representation of relation interactions, and therefore enables zero-shot generalization to any KG.
@@
In the absence of existing datasets for our inductive generalization setup, we curate a novel suite of 11 inductive query answering datasets where graphs and queries at inference time have new entity and relation vocabularies. Experimentally, we train a single UltraQuery model on one dataset and probe on other 22 transductive and inductive datasets. Averaged across the datasets, a single UltraQuery model outperforms by 50% (relative MRR) the best reported baselines in the literature (often tailored to specific graphs) on both EPFO queries and queries with negation.
@@ 2 Related Work
Due to the necessity of learning entity and relation embeddings, the vast majority of existing methods like GQE (Hamilton et al., 2018), BetaE (Ren & Leskovec, 2020), ConE (Zhang et al., 2021b), MPQE (Daza & Cochez, 2020) (and many more from the survey by Ren et al. (2023)) are transductive-only and tailored for a specific set of entities and relations. Among them, CQD (Arakelyan et al., 2021) and QTO (Bai et al., 2023) are inference-only query answering engines that execute logical operators with non-parametric fuzzy logic operators (e.g., product logic) but still require pre-trained entity and relation embedding matrices to execute relation projections (link prediction steps). We refer the interested reader to the comprehensive survey by Ren et al. (2023) that covers query answering theory, a taxonomy of approaches, datasets, and open challenges.
@@
A few models (Zhu et al., 2022; Galkin et al., 2022b; Gebhart et al., 2023) generalize only to new entities by modeling entities as a function of relation embeddings. Gebhart et al. (2023) apply the idea of cellular sheaves and harmonic extension to translation-based embedding models to answer conjunctive queries (without unions and negations). NodePiece-QE (Galkin et al., 2022b) trains an inductive entity encoder (based on the fixed vocabulary of relations) that is able to reconstruct entity embeddings of the new graph and then apply non-parametric engines like CQD to answer queries against new entities. The most effective inductive (entity) approach is GNN-QE (Zhu et al., 2022; Galkin et al., 2022b) that parameterizes each entity as a function of the relational structure between the query constants and the entity itself. However, all these works rely on a fixed relation vocabulary and cannot generalize to KGs with new relations at test time. In contrast, our model uses inductive relation projection and inductive logical operations that enable zero-shot generalization to any new KG with any entity and relation vocabulary without any specific training.
@@
Inductive Knowledge Graph Completion. In CLQA, KG completion methods execute the projection operator and are mainly responsible for predicting missing links in incomplete graphs during query execution. Inductive KG completion is usually categorized (Chen et al., 2023) into two branches: (i) inductive entity (inductive $(e)$) approaches have a fixed set of relations and only generalize to new entities, for example, to different subgraphs of one larger KG with one set of relations; and (ii) inductive entity and relation (inductive $(e,r)$) approaches that do not rely on any fixed set of entities and relations and generalize to any new KG with arbitrary new sets of entities and relations.
@@
Up until recently, the majority of existing approaches belonged to the inductive $(e)$ family (e.g., GraIL (Teru et al., 2020), NBFNet (Zhu et al., 2021), RED-GNN (Zhang & Yao, 2022), NodePiece (Galkin et al., 2022a), A*Net (Zhu et al., 2023), AdaProp (Zhang et al., 2023), inter alia) that generalizes only to new entities as their featurization strategies are based on learnable relation embeddings.
@@
Recently, the more generalizable inductive $(e,r)$ family started getting more attention, e.g., with RMPI (Geng et al., 2023), InGram (Lee et al., 2023), Ultra (Galkin et al., 2023), and the theory of double equivariance introduced by Gao et al. (2023) followed by ISDEA and MTDEA (Zhou et al., 2023) models. In this work, we employ Ultra to obtain transferable graph representations and execute the projection operator with link prediction over any arbitrary KG without input features. Extending our model with additional input features is possible (although deriving a single fixed-width model for graphs with arbitrary input space is highly non-trivial) and we leave it for future work.
@@ 3 Preliminaries and Problem Definition
We introduce the basic concepts pertaining to logical query answering and KGs largely following the formalizations introduced in Galkin et al. (2022b), Ren et al. (2023), and Galkin et al. (2023).
Knowledge Graphs and Inductive Setup. Given a finite set of entities $\mathcal{V}$ (nodes), a finite set of relations $\mathcal{R}$ (edge types), and a set of triples (edges) $\mathcal{E}=(\mathcal{V}\times\mathcal{R}\times\mathcal{V})$, a knowledge graph $\mathcal{G}$ is a tuple $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{E})$. The simplest transductive setup dictates that the graph at training time $\mathcal{G}_{\textit{train}}=(\mathcal{V}_{\textit{train}},\mathcal{R}_{\textit{train}},\mathcal{E}_{\textit{train}})$ and the graph at inference (validation or test) time $\mathcal{G}_{\textit{inf}}=(\mathcal{V}_{\textit{inf}},\mathcal{R}_{\textit{inf}},\mathcal{E}_{\textit{inf}})$ are the same, i.e., $\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}}$. By default, we assume that the inference graph $\mathcal{G}_{\textit{inf}}$ is an incomplete part of a larger, non observable graph $\hat{\mathcal{G}_{\textit{inf}}}$ with missing triples to be predicted at inference time. In the inductive setup, in the general case, the training and inference graphs are different, $\mathcal{G}_{\textit{train}}\neq\mathcal{G}_{\textit{inf}}$. In the easier inductive entity (inductive $(e)$) setup tackled by most of the KG completion literature, the relation set $\mathcal{R}$ is fixed and shared between training and inference graphs, i.e., $\mathcal{G}_{\textit{train}}=(\mathcal{V}_{\textit{train}},\mathcal{R},\mathcal{E}_{\textit{train}})$ and $\mathcal{G}_{\textit{inf}}=(\mathcal{V}_{\textit{inf}},\mathcal{R},\mathcal{E}_{\textit{inf}})$. The inference graph can be an extension of the training graph if $\mathcal{V}_{\textit{train}}\subseteq\mathcal{V}_{\textit{inf}}$ or be a separate disjoint graph (with the same set of relations) if $\mathcal{V}_{\textit{train}}\cap\mathcal{V}_{\textit{inf}}=\varnothing$. In CLQA, the former setup with the extended training graph at inference is tackled by InductiveQE approaches (Galkin et al., 2022b).
@@
In the inductive entity $(e)$ setup considered in Galkin et al. (2022b), the inference graph extends the training graph $\mathcal{G}_{\textit{train}}\subset\mathcal{G}_{\textit{inf}}$ but the set of relations is still fixed. Therefore, the proposed models are still bound to a certain hardcoded set of relations and cannot generalize to any arbitrary KG.
@@
Figure 2: (a) Example of ip query answering with UltraQuery: the inductive parametric projection operator (Section 4.1) executes relation projections on any graph and returns a scalar score for each entity; the scores are aggregated by non-parametric logical operators (Section 4.2) implemented with fuzzy logics. Intermediate scores are used for weighted initializion of relation projections on the next hop. (b) The multi-source propagation issue with a pre-trained link predictor for relation projection: pre-training on simple 1p link prediction is done in the single-source labeling mode (top) where only one query node is labeled with a non-zero vector; complex queries at later hops might have several plausible sources with non-zero initial weights from intermediate results (bottom) where a pre-trained operator fails.
Labeling Trick GNNs. Labeling tricks (as coined by Zhang et al. (2021a)) are featurization strategies in graphs for breaking symmetries in node representations which are particularly pronounced in link prediction and KG completion tasks. In the presence of such node symmetries (automorphisms), classical uni- and multi-relational GNN encoders (Kipf & Welling, 2017; Veličković et al., 2018; Vashishth et al., 2020) assign different automorphic nodes the same feature making them indistinguishable for downstream tasks. In multi-relational graphs, NBFNet (Zhu et al., 2021) and A*Net (Zhu et al., 2023) apply a labeling trick by using the indicator function $\textsc{Indicator}(h,v,r)$ that puts a query vector $\bm{r}$} on a head node $h$ and puts the zeros vector on other nodes $v$. The indicator function does not require entity embeddings and such models can naturally generalize to new entities (while the set of relation types is still fixed). Theoretically, such a labeling strategy learns conditional node representations and is provably more powerful (Huang et al., 2023) than node-level GNN encoders. In CLQA, only GNN-QE (Zhu et al., 2022) applies NBFNet as a projection operator making it the only approach generalizable to the inductive $(e)$ setup (Galkin et al., 2022b) with new nodes at inference time. This work leverages labeling trick GNNs to generalize CLQA to arbitrary KGs with any entity and relation vocabulary.
@@ 4 Method
We aim at designing a single model for CLQA on any KG in the zero-shot fashion, i.e., without training on a target graph. In the CLQA literature (Hamilton et al., 2018; Ren et al., 2020; Ren & Leskovec, 2020; Arakelyan et al., 2021; Zhu et al., 2022), it is common to break down query execution into a relation projection to traverse graph edges and predict missing links, and logical operators that model conjunction, disjunction, and union. Therefore, the main challenge boils down to designing inductive projection and logical operators suitable for any (even new) entity and relation vocabulary.

### 01 Oct 2024 v2

# A Foundation Model for Zero-shot Logical Query Reasoning

@@ Abstract
Complex logical query answering (CLQA) in knowledge graphs (KGs) goes beyond simple KG completion and aims at answering compositional queries comprised of multiple projections and logical operations. Existing CLQA methods that learn parameters bound to certain entity or relation vocabularies can only be applied to the graph they are trained on which requires substantial training time before being deployed on a new graph. Here we present UltraQuery, the first foundation model for inductive reasoning that can zero-shot answer logical queries on any KG. The core idea of UltraQuery is to derive both projections and logical operations as vocabulary-independent functions which generalize to new entities and relations in any KG. With the projection operation initialized from a pre-trained inductive KG completion model, UltraQuery can solve CLQA on any KG after finetuning on a single dataset. Experimenting on 23 datasets, UltraQuery in the zero-shot inference mode shows competitive or better query answering performance than best available baselines and sets a new state of the art on 15 of them.
Figure 1: Zero-shot query answering performance (MRR, higher is better) of a single UltraQuery model trained on one FB15k237 queries dataset compared to the best available baselines and ablated UltraQuery LP on 23 datasets. EPFO is the average of 9 query types with $(\wedge,\lor)$ operators, Negation is the average of 5 query types with the negation operator $(\neg)$. On average, a single UltraQuery model outperforms the best baselines trained specifically on each dataset. More results are presented in Table 2 and Appendix C.
@@ 1 Introduction
Complex logical query answering (CLQA) generalizes simple knowledge graph (KG) completion to more complex, compositional queries with logical operators such as intersection $(\wedge)$, union $(\lor)$, and negation $(\lnot)$. Such queries are expressed in a subset of first-order logic (FOL) where existentially quantified $(\exists)$ variables and given constants comprise relation projections (or atoms), and logical operators combine projections into a logical query (graph pattern). A typical example of a logical query [27] is presented in Figure 2: $U.\exists V:\texttt{Win}(\texttt{NobelPrize},V)\land\texttt{Citizen}(\texttt{USA},V)\land\texttt{Graduate}(V,U)$ where $\texttt{Win}()$ is a relation projection, NobelPrize is a constant, and $V$ is an existentially quantified variable.
@@
Due to the incompleteness of most KGs, these logical queries cannot be directly solved by graph traversal algorithms. Consequently, CLQA methods have to deal with missing edges when modeling the projection operators. The vast majority of existing CLQA methods [27, 26, 25, 3, 5] predict missing edges by learning graph-specific entity and relation embeddings making such approaches transductive and not generalizable to other KGs. A few approaches [43, 14, 17] are able to generalize query answering to new nodes at inference time but still need a fixed relation vocabulary.
@@
In this work, we focus on the hardest inductive generalization setup where queries and underlying graphs at inference time are completely different from the training graph, i.e., both entities and relations are new. Furthermore, we aim at performing CLQA in the zero-shot setting with one single model. That is, instead of finetuning a model on each target dataset, we seek to design a unified approach that generalizes to any KG and query at inference time. For example, in Figure 2, the training graph describes academic entities with relations Win, Citizen, Graduate [^1] We assume the presence of respective inverse relations $\texttt{r}^{-1}$. whereas the inference graph describes music entities with relations Band Member and Plays. The query against the inference graph $U:\texttt{BandMember}(\texttt{Dire Straits},U)\wedge\texttt{Plays}^{-1}(\texttt{Guitar},U)$ involves both new entities and relations and, to the best of our knowledge, cannot be tackled by any existing CLQA method that learns a fixed set of entities or relation embeddings from the training graph.
@@
Contributions. Our contributions are two-fold. First, none of the existing CLQA methods can generalize to query answering over new arbitrary KGs with new entities and relations at inference time. We bridge this gap by leveraging the recent progress in inductive KG reasoning [15, 16] and devise UltraQuery, the first foundation model for CLQA that generalizes to logical queries on any arbitrary KG with any entity and relation vocabulary in the zero-shot fashion without relying on any external node or edge features. UltraQuery parameterizes the projection operator by an inductive graph neural network (GNN) and implements non-parametric logical operators with fuzzy logics [31]. The pre-trained projection operator [15] does not learn any graph-specific entity nor relation embeddings thanks to the generalizable meta-graph representation of relation interactions, and therefore enables zero-shot generalization to any KG.
@@
Second, in the absence of existing datasets for our inductive generalization setup, we curate a novel suite of 11 inductive query answering datasets where graphs and queries at inference time have new entity and relation vocabularies. Experimentally, we train a single UltraQuery model on one dataset and probe on other 22 transductive and inductive datasets. Averaged across the datasets, a single UltraQuery model outperforms by 50% (relative MRR) the best reported baselines in the literature (often tailored to specific graphs) on both EPFO queries and queries with negation.
@@ 2 Related Work
Due to the necessity of learning entity and relation embeddings, the vast majority of existing methods like GQE [19], BetaE [25], ConE [40], MPQE [10] (and many more from the survey by Ren et al. [27]) are transductive-only and tailored for a specific set of entities and relations. Among them, CQD [3] and QTO [5] are inference-only query answering engines that execute logical operators with non-parametric fuzzy logic operators (e.g., product logic) but still require pre-trained entity and relation embedding matrices to execute relation projections (link prediction steps). We refer the interested reader to the comprehensive survey by Ren et al. [27] that covers query answering theory, a taxonomy of approaches, datasets, and open challenges.
Table 1: Comparison with existing CLQA approaches. Ind. denotes inductive generalization to new entities (e) and relations (r). UltraQuery is the first inductive method the generalizes to queries over new entities and relations at inference time.
[…]
@@
A few models [43, 14, 17] generalize only to new entities by modeling entities as a function of relation embeddings. Gebhart et al. [17] apply the idea of cellular sheaves and harmonic extension to translation-based embedding models to answer conjunctive queries (without unions and negations). NodePiece-QE [14] trains an inductive entity encoder (based on the fixed vocabulary of relations) that is able to reconstruct entity embeddings of the new graph and then apply non-parametric engines like CQD to answer queries against new entities. The most effective inductive (entity) approach is GNN-QE [43, 14] that parameterizes each entity as a function of the relational structure between the query constants and the entity itself. However, all these works rely on a fixed relation vocabulary and cannot generalize to KGs with new relations at test time. In contrast, our model uses inductive relation projection and inductive logical operations that enable zero-shot generalization to any new KG with any entity and relation vocabulary without any specific training.
@@
Inductive Knowledge Graph Completion. In CLQA, KG completion methods execute the projection operator and are mainly responsible for predicting missing links in incomplete graphs during query execution. Inductive KG completion is usually categorized [8] into two branches: (i) inductive entity (inductive $(e)$) approaches have a fixed set of relations and only generalize to new entities, for example, to different subgraphs of one larger KG with one set of relations; and (ii) inductive entity and relation (inductive $(e,r)$) approaches that do not rely on any fixed set of entities and relations and generalize to any new KG with arbitrary new sets of entities and relations.
@@
Up until recently, the majority of existing approaches belonged to the inductive $(e)$ family (e.g., GraIL [29], NBFNet [42], RED-GNN [38], NodePiece [13], A*Net [44], AdaProp [39]) that generalizes only to new entities as their featurization strategies are based on learnable relation embeddings.
@@
Recently, the more generalizable inductive $(e,r)$ family started getting more attention, e.g., with RMPI [18], InGram [22], Ultra [15], and the theory of double equivariance introduced by Gao et al. [16] followed by ISDEA and MTDEA [41] models. In this work, we employ Ultra to obtain transferable graph representations and execute the projection operator with link prediction over any arbitrary KG without input features. Extending our model with additional input features is possible (although deriving a single fixed-width model for graphs with arbitrary input space is highly non-trivial) and we leave it for future work.
@@ 3 Preliminaries and Problem Definition
We introduce the basic concepts pertaining to logical query answering and KGs largely following the existing literature [14, 27, 15].
Knowledge Graphs and Inductive Setup. Given a finite set of entities $\mathcal{V}$ (nodes), a finite set of relations $\mathcal{R}$ (edge types), and a set of triples (edges) $\mathcal{E}=(\mathcal{V}\times\mathcal{R}\times\mathcal{V})$, a knowledge graph $\mathcal{G}$ is a tuple $\mathcal{G}=(\mathcal{V},\mathcal{R},\mathcal{E})$. The simplest transductive setup dictates that the graph at training time $\mathcal{G}_{\textit{train}}=(\mathcal{V}_{\textit{train}},\mathcal{R}_{\textit{train}},\mathcal{E}_{\textit{train}})$ and the graph at inference (validation or test) time $\mathcal{G}_{\textit{inf}}=(\mathcal{V}_{\textit{inf}},\mathcal{R}_{\textit{inf}},\mathcal{E}_{\textit{inf}})$ are the same, i.e., $\mathcal{G}_{\textit{train}}=\mathcal{G}_{\textit{inf}}$. By default, we assume that the inference graph $\mathcal{G}_{\textit{inf}}$ is an incomplete part of a larger, non observable graph $\hat{\mathcal{G}_{\textit{inf}}}$ with missing triples to be predicted at inference time. In the inductive setup, in the general case, the training and inference graphs are different, $\mathcal{G}_{\textit{train}}\neq\mathcal{G}_{\textit{inf}}$. In the easier inductive entity (inductive $(e)$) setup tackled by most of the KG completion literature, the relation set $\mathcal{R}$ is fixed and shared between training and inference graphs, i.e., $\mathcal{G}_{\textit{train}}=(\mathcal{V}_{\textit{train}},\mathcal{R},\mathcal{E}_{\textit{train}})$ and $\mathcal{G}_{\textit{inf}}=(\mathcal{V}_{\textit{inf}},\mathcal{R},\mathcal{E}_{\textit{inf}})$. The inference graph can be an extension of the training graph if $\mathcal{V}_{\textit{train}}\subseteq\mathcal{V}_{\textit{inf}}$ or be a separate disjoint graph (with the same set of relations) if $\mathcal{V}_{\textit{train}}\cap\mathcal{V}_{\textit{inf}}=\varnothing$. In CLQA, the former setup with the extended training graph at inference is tackled by InductiveQE approaches [14].
@@
In the inductive entity $(e)$ setup considered in Galkin et al. [14], the inference graph extends the training graph $\mathcal{G}_{\textit{train}}\subset\mathcal{G}_{\textit{inf}}$ but the set of relations is still fixed. Therefore, the proposed models are still bound to a certain hardcoded set of relations and cannot generalize to any arbitrary KG.
@@
Labeling Trick GNNs. Labeling tricks (as coined by Zhang et al. [37]) are featurization strategies in graphs for breaking symmetries in node representations which are particularly pronounced in link prediction and KG completion tasks. In the presence of such node symmetries (automorphisms), classical uni- and multi-relational GNN encoders [21, 33, 32] assign different automorphic nodes the same feature making them indistinguishable for downstream tasks. In multi-relational graphs, NBFNet [42] and A*Net [44] apply a labeling trick by using the indicator function $\textsc{Indicator}(h,v,r)$ that puts a query vector $\bm{r}$} on a head node $h$ and puts the zeros vector on other nodes $v$. The indicator function does not require entity embeddings and such models can naturally generalize to new entities (while the set of relation types is still fixed). Theoretically, such a labeling strategy learns conditional node representations and is provably more powerful [20] than node-level GNN encoders. In CLQA, only GNN-QE [43] applies NBFNet as a projection operator making it the only approach generalizable to the inductive $(e)$ setup [14] with new nodes at inference time. This work leverages labeling trick GNNs to generalize CLQA to arbitrary KGs with any entity and relation vocabulary.
Figure 3: (a) Example of ip query answering with UltraQuery: the inductive parametric projection operator (Section 4.1) executes relation projections on any graph and returns a scalar score for each entity; the scores are aggregated by non-parametric logical operators (Section 4.2) implemented with fuzzy logics. Intermediate scores are used for weighted initializion of relation projections on the next hop. (b) The multi-source propagation issue with a pre-trained link predictor for relation projection: pre-training on 1p link prediction is done in the single-source labeling mode (top) where only one query node is labeled with a non-zero vector; complex queries at later intermediate hops might have several plausible sources with non-zero initial weights (bottom) where a pre-trained operator fails.
@@ 4 Method
We aim at designing a single foundation model for CLQA on any KG in the zero-shot fashion, i.e., without training on a target graph. In the CLQA literature [19, 26, 25, 3, 43], it is common to break down query execution into a relation projection to traverse graph edges and predict missing links, and logical operators that model conjunction, disjunction, and union. The main challenge boils down to designing inductive projection and logical operators suitable for any entity and relation vocabulary.